I need to find the lenght of a hypotenuse (the rope). A flagpole is 45' high. A guy wire is tied from the top to make a 45 degree angle w/the ground. what is the length of the rope? so i know i have 90, 45 & 45 for my angles, & the height is 45'. HELP!! Thanks (this is for my 5th grade son).
Answer by venugopalramana(3286)

I need to find the lenght of a hypotenuse (the rope). A flagpole is 45' high. A guy wire is tied from the top to make a 45 degree angle w/the ground. what is the length of the rope? so i know i have 90, 45 & 45 for my angles, & the height is 45'....CORRECT.IN A TRIANGLE IF 2 ANGLES ARE SAME THEIR OPPOSITE SIDES ARE ALSO EQUAL
HENCE THE 2 SIDES ARE EQUAL TO 45'.
HYPOTENUSE^2=SIDE^2+SIDE^2=45^2+45^2=2025+2025=4050
HYPOTENUSE=SQRT(4050)=63.64
HELP!! Thanks (this is for my 5th grade son).